appl112:

What is the value of r in the equation 2(2r + 8) = 3r + 16?

silvernight269:

First you have to distribute the 2. So 2 times 2r and 2 times 8 to equal 4r+16. Then 4r -3r and 16-16 to equal r=0

Symphony:

\(4r+16=3r+16\) \(4r-3r=16-16\) \(r=0\)